# PooDLe[Uncaptioned image]: Pooled and dense self-supervised learning from naturalistic videos

### v2

@@ 1 Introduction
    We introduce PooDLe, a new SSL method which overcomes the challenges of spatial imbalance and cluttered scenes by unifying a flow equivariance, dense SSL objective and a pooled objective over pseudo-iconic subcrops alongside a spatial decoder module to effectively learn from naturalistic video.
@@
    We study the effects of global and subcrop area, input resolution and temporal stride between paired frames.
@@ 4.2 Main Results
PooDLe also outperforms supervised IN1K pretraining, despite the latter’s advantage in learning small and rare classes present in BDD100K (spatial imbalance shown in Figure 1) due to ImageNet being a class-balanced dataset with iconic views of objects. In addition, we pretrain PooDLe on BDD100K with weights initialized from the supervised IN1K checkpoint, improving mIoU by $8\%$ and Acc by $6\%$ over the initialization weights on linear semantic segmentation. In Appendix F, we show PooDLe remains competitive against IN1K-pretrained baselines despite being trained in the challenging naturalistic video setting.
@@
We also train PooDLe on $\text{WT}_{\text{Venice}}$ and $\text{WT}_{\text{all}}$. Table 2 shows results on ADE20K (Zhou et al., 2017) semantic segmentation using linear readout and finetuning following (Venkataramanan et al., 2024). Notably, when pretrained on $\text{WT}_{\text{all}}$, PooDLe obtains $2.6\%$ higher mIoU than DoRA on ADE20K linear readout, and $2.7\%$ mIoU on UperNet finetuning. PooDLe also performs better on $\text{WT}_{\text{Venice}}$, with a gain of $1.4\%$ mIoU over DoRA and $0.6\%$ mIoU over PixPro on ADE20K UperNet finetuning. Figure 16 shows predicted segmentation masks for ADE20K.
@@
Naturalistic videos have imbalanced class representation and object sizes (Figure 1; e.g., “road” occupies $21\%$ of pixels while “bicycle” only occupies $0.05\%$ of pixels). Capturing information on these underrepresented classes is very challenging. To further demonstrate this phenomenon, we categorize BDD classes as “small” if they occupy $<1\%$ of pixels and “large” for those that occupy $>1\%$. Separately, we define “rare” as classes that appear in $<20\%$ of images and “common” as those that appear in $>20\%$. Table 3 shows linear readout mIoU for different class groupings, highlighting the impact of class and spatial imbalance. Full class-level statistics and designations are in Appendix G. We observe that FlowE performs well on large and common classes due to its dense loss, but struggles on small or rare classes. Meanwhile, supervised IN1K, benefiting from balanced pretraining data, effectively learns about smaller classes. PooDLe, with its unified objectives and spatial decoder module, significantly outperforms other BDD-pretrained models across all class groupings, particularly on small and rare classes. PooDLe, initialized from supervised IN1K weights, significantly improves upon supervised IN1K on large classes, from $42.2\%$ to $56.1\%$, due to the dense objective, while remaining competitive with supervised IN1K on small classes.
@@ 4.3 Ablation studies
We also demonstrate that PooDLe is able to perform well even with self-supervised flow. We train UFlow (Jonschkowski et al., 2020) model on KITTI (Geiger et al., 2013) and finetune it on BDD resulting in only a 0.5% mIoU loss compared to pretraining with RAFT. See Appendix A and I for details and visualizations.
@@ 4.4 Spatial and temporal cropping in self-supervised video learning
Figure 9:
@@
We select 3 different resolutions, and sample the crop area from a Gaussian truncated to $[0,1]$, with varying mean and $\sigma=0.1$. All three resolutions are trained with $\mu=0.125,0.275,0.425,0.725$ with the 2 smaller resolutions also trained at $\mu=0.05,0.20$ for higher granularity. The largest model ($512\times 1024$) produces the best results and peaks at $\mu=0.425$ while the other two peak at smaller crop areas. The $256\times 512$ model degrades more slowly in performance as crop area increases in comparison to the $128\times 256$ model.
@@
When it is small, there is limited variance in object appearance, diminishing the value of video data, and when it is too large, correspondence between frames decreases and optical flow becomes unreliable. Note that for $\Delta t=0$, we add jitter to the initial large crop by up to $10\%$ of the image size. Figure 9 shows frame sequences from 2 different videos, highlighting the high variability of motion in BDD100K.
@@ Appendix E Additional visualizations
Appendix F Additional evaluation results
@@
Appendix G Per-class evaluation results
@@
Appendix H Backbone computation cost
@@
Appendix I Flow visualizations

### v3

@@ 1 Introduction
    We introduce PooDLe, a new SSL method which overcomes the challenges of spatial imbalance and cluttered scenes by unifying a flow equivariance, dense SSL objective and a pooled objective over pseudo-iconic subcrops alongside a spatial decoder module to effectively learn from naturalistic video. PooDLe achieves state-of-the-art performance on BDD100K (Yu et al., 2020) and Cityscapes (Cordts et al., 2016) semantic segmentation and BDD object detection. It also obtains the highest mIoU on ADE20K (Zhou et al., 2017), and on WT-Sem, our new, in-distribution semantic segmentation task for Walking Tours.
@@
    We study the effects of global and subcrop area, input resolution and temporal stride between paired frames. We show the importance of maintaining pixel density by adjusting crop area when training with larger resolutions for learning visual representations. We also verified that smaller subcrop areas are able to better capture smaller classes. We believe these observations will be helpful in guiding future work on dense, naturalistic data.
@@ 4.2 Main Results
PooDLe also outperforms supervised IN1K pretraining, despite the latter’s advantage in learning small and rare classes present in BDD100K (spatial imbalance shown in Figure 1) due to ImageNet being a class-balanced dataset with iconic views of objects. In addition, we pretrain PooDLe on BDD100K with weights initialized from the supervised IN1K checkpoint, improving mIoU by $8\%$ and Acc by $6\%$ over the initialization weights on linear semantic segmentation. In Appendix G, we show PooDLe remains competitive against IN1K-pretrained baselines despite being trained in the challenging naturalistic video setting.
@@
We also train PooDLe on $\text{WT}_{\text{Venice}}$ and $\text{WT}_{\text{all}}$. Table 2 shows results on ADE20K (Zhou et al., 2017) semantic segmentation using linear readout and finetuning following (Venkataramanan et al., 2024). Notably, when pretrained on $\text{WT}_{\text{all}}$, PooDLe obtains $2.6\%$ higher mIoU than DoRA on ADE20K linear readout, and $2.7\%$ mIoU on UperNet finetuning. PooDLe also performs better on $\text{WT}_{\text{Venice}}$, with a gain of $1.4\%$ mIoU over DoRA and $0.6\%$ mIoU over PixPro on ADE20K UperNet finetuning. We note that PooDLe uses a smaller ResNet-50 backbone and is trained for fewer epochs than DoRA, the strongest baseline. Despite these differences, these results show that PooDLe learns strong representations from naturalistic video captured in the open world. Figure 16 shows predicted segmentation masks for ADE20K.
@@
Naturalistic videos have imbalanced class representation and object sizes (Figure 1; e.g., “road” occupies $21\%$ of pixels while “bicycle” only occupies $0.05\%$ of pixels). Capturing information on these underrepresented classes is very challenging. To further demonstrate this phenomenon, we categorize BDD classes as “small” if they occupy $<1\%$ of pixels and “large” for those that occupy $>1\%$. Separately, we define “rare” as classes that appear in $<20\%$ of images and “common” as those that appear in $>20\%$. Table 3 shows linear readout mIoU for different class groupings, highlighting the impact of class and spatial imbalance. Full class-level statistics and designations are in Appendix H. We observe that FlowE performs well on large and common classes due to its dense loss, but struggles on small or rare classes. Meanwhile, supervised IN1K, benefiting from balanced pretraining data, effectively learns about smaller classes. PooDLe, with its unified objectives and spatial decoder module, significantly outperforms other BDD-pretrained models across all class groupings, particularly on small and rare classes. PooDLe, initialized from supervised IN1K weights, significantly improves upon supervised IN1K on large classes, from $42.2\%$ to $56.1\%$, due to the dense objective, while remaining competitive with supervised IN1K on small classes.
@@ 4.3 Ablation studies
We also demonstrate that PooDLe is able to perform well even with self-supervised flow. We train UFlow (Jonschkowski et al., 2020) model on KITTI (Geiger et al., 2013) and finetune it on BDD resulting in only a 0.5% mIoU loss compared to pretraining with RAFT. See Appendix A and J for details and visualizations.
@@ 4.4 Spatial and temporal cropping in self-supervised video learning
Figure 9: mIoU when varying $\Delta t$.
@@
Next, we vary both the global crop area from the raw video frame along with the input resolution to study their effects on self-supervised pretraining. We select 3 different resolutions, and sample the crop area from a Gaussian truncated to $[0,1]$, with varying mean and $\sigma=0.1$. All three resolutions are trained with $\mu=0.125,0.275,0.425,0.725$ with the 2 smaller resolutions also trained at $\mu=0.05,0.20$ for higher granularity. Our results in Figure 5 show that larger crop areas and higher input resolutions, together, are important for maximizing performance. The largest model ($512\times 1024$) produces the best results and peaks at $\mu=0.425$ while the other two peak at smaller crop areas. The $256\times 512$ model degrades more slowly in performance as crop area increases in comparison to the $128\times 256$ model.
@@
We study the effect of temporal stride by training PooDLe with $\Delta t=0,8,15,30,45$ on BDD100K, evaluated using linear readout (Figure 9). Performance peaks at $\Delta t=15$ and degrades only slightly at $8$ and, $30$ while dropping further at values 00 and $45$. When it is small, there is limited variance in object appearance, diminishing the value of video data, and when it is too large, correspondence between frames decreases and optical flow becomes unreliable. Note that for $\Delta t=0$, we add jitter to the initial large crop by up to $10\%$ of the image size. Figure 9 shows frame sequences from 2 different videos, highlighting the high variability of motion in BDD100K.
@@ Appendix E Additional visualizations
Appendix F Accuracy values for class breakdown and ablations
Table 6: Accuracy values for Table 3, class breakdowns
Method Pretrain All Small Large Rare Common
DINO BDD 86.8 12.3 88.3 2.3 87.8
DenseCL BDD 84.9 2.0 86.6 0.0 86.0
DoRA BDD 88.1 19.3 89.5 7.2 89.1
FlowE BDD 88.5 18.2 89.9 32.0 89.2
PooDLe BDD 89.2 33.6 90.3 34.2 89.9
Supervised IN1K 84.7 36.9 85.3 23.8 85.1
PooDLe BDD* 90.7 35.6 91.2 46.9 91.2
Table 7: Accuray values for Table 4, ablations
Variant Dense Pool Top-Down Lateral Flow All Small Large Rare Common
1 FlowE RAFT 85.0 22.8 86.3 6.1 86.0
2 RAFT 86.2 14.2 87.6 6.3 87.1
3 RAFT 86.8 11.9 87.7 13.6 87.7
4 RAFT 86.6 22.1 87.9 16.9 87.5
5 RAFT 84.2 25.5 85.3 28.2 84.9
[…]
7 PooDLe RAFT 86.5 26.6 87.7 28.5 87.1
Appendix G Additional evaluation results
@@
Appendix H Per-class evaluation results
@@
Appendix I Backbone computation cost
@@
Appendix J Flow visualizations